OpenStudy (karolynxx):

Which value, when placed in the box, would result in a system of equations with infinitely many solutions? y = 2x – 5 2y – 4x = a. –10 b. –5 c. 5 d. 10

OpenStudy (karolynxx):

@sleepyjess do you know what i have to do?

OpenStudy (sleepyjess):

First, let's get these into the same format. I would probably subtract 2x from both sides of the first equation.

OpenStudy (karolynxx):

so 2x-2y?

OpenStudy (sleepyjess):

Close, y - 2x = -5 So, y - 2x = -5 2y - 4x = ? What do you notice between the x's and y's in each equation?

OpenStudy (karolynxx):

their all lined up in the same place

OpenStudy (karolynxx):

i got -10 and that was the anwser (: thank you!

OpenStudy (sleepyjess):

Yes, we did that to see the correlation between them :)

OpenStudy (sleepyjess):

Good! The second equation was double the first equation :3